OpenStudy (anonymous):

Select the number property that will justify the step in the following expression. 6 + (14 + 70) = (6 + 14) + 70 = Identity-Addition, Transitive, Commutative-Addition, Associative-Addition 20 + 70 = 90

OpenStudy (anonymous):

Associative addition should google each of them to see some examples

OpenStudy (mathstudent55):

Identity-Addition: a + 0 = a Transitive Property: If a = b, and b = c, then a = c Commutative-Addition: a + b = b + a Associative-Addition: a + (b + c) = (a + b) + c Can you choose now?

OpenStudy (anonymous):

i agree to get (6 + 14) + 70 = from 6 + (14 + 70) you must used associative addition